OpenStudy (anonymous):

please find the value of c that makes the expression a perfect square trinomial, then write the expression as a square of a binomial x^2+12x+c

OpenStudy (campbell_st):

for a perfect square (x +a)^2 the expansion is x^2 + 2ax + a^2 in your question 2a = 12 so a = 6 a^2 = (6)^2 = 36 for the perfect square in your question c must equal 36